OpenStudy (anonymous):

please help 1 medal and fan Find the slope of a line perpendicular to 2x-y=16 . 2 - 1/2 8 1/2

OpenStudy (espex):

Put your equation into slope-intercept (\(y=mx+b\)) form.

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

- 1/2

OpenStudy (espex):

Isolating y, to get the slope-intercept equation: -y = 16 - 2x y = -(16 - 2x) y = 2x - 16 y = mx + b

OpenStudy (anonymous):

2

OpenStudy (espex):

Exactly. :)